OpenStudy (anonymous):

Need to separate and integrate dv/dt = v+2t

OpenStudy (turingtest):

This differential equation does not look separable to me. Are you sure you typed it correctly?

OpenStudy (anonymous):

This equation is not separable. Rather, the easiest way to solve it is with an integrating factor.

OpenStudy (turingtest):

v'-v=2t integrating factor:e^t (ve^t)'=2te^tdt now integrate

OpenStudy (jamesj):

No, integrating factor is e^(-t)

OpenStudy (turingtest):

my bad...

OpenStudy (jamesj):

MIT should pay me a commission ... but @jd50, if this method is obscure, watch this solid lecture: http://ocw.mit.edu/courses/mathematics/18-03-differential-equations-spring-2010/video-lectures/lecture-3-solving-first-order-linear-odes/

OpenStudy (turingtest):

v'-v=2t multiply by the integrating factor:e^(-t) [(ve^(-t)]'=2te^(-t)dt now integrate

OpenStudy (turingtest):

I learned my technique from this site: http://tutorial.math.lamar.edu/Classes/DE/Linear.aspx

myininaya (myininaya):

\[v'-v=2t\] multiply both sides by y(x)>0 such that we have \[yv'-yv=y \cdot 2t\] we want to write the right hand side as (vy)' to do that we need y'=-y => dy/dt=-y =>-1/y dy=dt integrating both sides we get -ln(y)=t+C we choose C=0 so we have t=-ln(y) => -t=ln(y)=> e^(-t)=y so we can write \[(e^{-t}v)'=e^{-t} \cdot 2t\]

myininaya (myininaya):

oops multiply both sides by y(t) lol